# ACTISM: Threat-Informed Dynamic Security Modelling for Automotive Systems

### v2

@@ Abstract
Cybersecurity threats in complex cyber-physical systems pose significant risks to system functionality and safety. This experience report introduces ACTISM (Automotive Consequence-Driven and Threat-Informed Security Modelling) — an integrated security modeling approach that enhances the resilience of automotive systems by dynamically updating their cybersecurity posture in response to prevailing threats, attacker tactics, and their impact on system functionality and safety. ACTISM addresses the existing knowledge gap in static security assessment methodologies by providing a dynamic and iterative framework. We demonstrate the effectiveness of ACTISM by applying it to a real-world example of the Tesla Electric Vehicle’s In-Vehicle Infotainment system, illustrating how the security model can be adapted as new threats emerge. We also outline avenues for future research and development in this area, including automated vulnerability management workflows for automotive systems.
@@ 1 Introduction
Automotive systems have transformed into complex cyber-physical systems brimming with cutting-edge features. In contrast to their predecessors, today’s automotive systems are equipped with a variety of advanced sensors and connectivity options, such as Control Area Networks (CAN), Local Interconnect Networks (LIN), FlexRay, and Ethernet. They also offer external connectivity through Bluetooth, Wi-Fi, and GPS. In addition, the integration of feature-rich In-Vehicle Infotainment (IVI) systems and powerful computing units for autonomous driving continues to push the boundaries of what is possible in automotive technology. However, this technological evolution also broadens the attack surface, making automotive systems more susceptible to cybersecurity attacks.
Cybersecurity threats and vulnerabilities in the automotive industry are well documented, and attackers exploit these vulnerabilities to compromise vehicle security, access sensitive data, and even take control of vehicle functions.
Recent literature has increasingly highlighted the challenges of defending against cybersecurity threats in automotive systems. In their survey on security attacks and defence techniques for Connected and Autonomous Vehicles (CAVs), Pham and Xiong [11] observed that attackers could potentially compromise one vehicle component after another. To mitigate this risk, they emphasised the need for manufacturers to strengthen the security of all system components to ensure the overall protection of CAVs.
While this approach may be feasible during the early stages of an automotive system’s life cycle, evolving cybersecurity threats are likely to render it less effective over time. Given the rapid increase in both threats and the complexity of automotive systems, it is unrealistic to expect that an automotive system’s initial cybersecurity features will counter all emerging risks throughout its entire life cycle. This is where threat-informed dynamic threat and attack models become crucial, offering a proactive and adaptive approach to designing secure automotive systems.
Dynamic threat and attack models provide engineers with a structured framework to analyse and mitigate cybersecurity risks, ensuring that automotive systems remain protected against evolving cybersecurity threats. Additionally, they enable consumers to make informed decisions when selecting vehicles based on their cybersecurity robustness.
However, many of the threat and attack models discussed in the literature do not fully address the unique cybersecurity requirements of modern automotive systems. These requirements include:
@@
For instance, threat models developed using Microsoft’s STRIDE (Spoofing, Tampering, Repudiation, Information Disclosure, Denial of Service, and Elevation of Privilege) primarily focus on data flows and information security. However, STRIDE was not designed with automotive systems in mind, making its application in this context somewhat limited. Consequently, STRIDE often overlooks critical aspects of how cybersecurity breaches can directly affect automotive functions and safety. As a result, the reliability of the general outcomes from STRIDE is questionable when applied to the unique characteristics of automotive environments.
In other words, threat modelling of automotive systems often does not consider the consequences of cybersecurity attacks beyond information security. Similarly, attack models designed for traditional IT systems do not fully align with the multi-path tactics and techniques employed by attackers seeking to gain control of a vehicle. These techniques often involve a combination of physical and cyber actions, necessitating a more comprehensive and dynamic approach to security modelling.
@@
These findings highlight the need to tailor security modeling approaches for automotive systems so that they can adequately capture the unique cybersecurity threats faced in these environments.
@@
Contributions. In this experience report, we introduce ACTISM (Automotive Consequence-Driven and Threat-Informed Security Modelling) — an integrated security modeling approach that enhances the resilience of automotive systems by dynamically updating their cybersecurity posture in response to prevailing threats, attacker tactics, and their impact on system functionality and safety. This approach addresses the existing knowledge gap in static security assessment methodologies by providing a dynamic and iterative framework. We demonstrate the effectiveness of ACTISM by applying it to an IVI system, illustrating how the model can be adapted as new threats emerge.
@@
Unlike existing security modelling approaches, ACTISM integrates various frameworks to dynamically update security models with new information, acquired through threat intelligence and ongoing security monitoring. This keeps security models relevant against evolving attacker tactics, techniques and procedures (TTPs) even if the models were developed in early stages of system development. ACTISM also allows practitioners to apply the security models in both IT and automotive domains by using attack paths to capture the multi-path and multi-agent aspects of real-world cyber-physical attacks. In order to bolster the completeness of the security model, ACTISM integrates the following threat analysis and risk assessment frameworks:
  • Microsoft’s STRIDE methodology.
@@
  • HEAVENS (HEAling Vulnerabilities to ENhance Software Security and Safety), a structured and model-based hazard analysis and risk assessment method specifically designed for automotive systems.
@@
  • Attack Tree, a hierarchical diagram representing potential attack paths, illustrating how an attacker could exploit system vulnerabilities to achieve specific goals.
@@
  • Common Vulnerability Scoring System (CVSS), an open industry standard for assessing the severity of computer system security vulnerabilities. CVSS (Common Vulnerability Scoring System): This model is also static as it provides a standardised method for evaluating the severity of vulnerabilities at a particular point in time. It does not dynamically adapt to new threats without manual intervention.
@@
  • Threat Analysis and Risk Assessment (TARA) methodology from the ISO/SAE 21434:2021 ("Road Vehicles - Cybersecurity Engineering") Standard, jointly developed by the International Organisation for Standardisation (ISO) and the Society of Automotive Engineers (SAE), for managing cybersecurity threats and vulnerabilities throughout the life cycle of a vehicle.
Table 1 lists the differences between ACTISM and these security modelling frameworks, and their respective limitations. In contrast with other security modelling frameworks, the ACTISM framework integrates various frameworks to be dynamic, continuously adapting to new threats and updating the automotive system’s cybersecurity posture. This distinction sets it apart from the static models, providing a more robust approach to maintaining automotive cybersecurity.
Table 1: Comparison of ACTISM with other security modelling frameworks
Framework Type Limitations
ACTISM Dynamic Relies on the integration of various security modelling frameworks.
STRIDE Static Limited to specific threat categories. Does not dynamically adapt to new threats without manual intervention.
Attack Trees Static Requires manual updates and may not cover all emerging threats. Does not dynamically adapt to new threats without manual intervention.
HEAVENS Static Primarily focuses on high-level security analysis which may not cover detailed, system-specific vulnerabilities and threats. Does not dynamically adapt to new threats without manual intervention.
CVSS Static Does not dynamically adapt to new threats without manual intervention.
TARA Static Does not dynamically adapt to new threats without manual intervention.
@@
Although frameworks such as STRIDE and CVSS were designed for IT security threats, they can also be applied to physical threats. As an example, tampering with physical components can lead to unauthorised access, data manipulation, or service disruption. As an integrative approach, ACTISM provides guidance on how both cyber and physical attacks can be modeled using these frameworks to capture the various entry points and methods attackers might use, ensuring comprehensive security coverage across both cyber and physical domains.
To demonstrate the practical application of ACTISM, we developed a comprehensive security model for an IVI system. We outline the step-by-step process of building the security model, emphasising the integration of threat intelligence and contextual information to accurately characterise and assess cybersecurity risks.
Furthermore, we demonstrate the utility of ACTISM by illustrating its adaptive response to a real-world vulnerability disclosure affecting the IVI system. This case study underscored ACTISM’s effectiveness in addressing emerging threats and evolving security challenges in automotive cybersecurity. Furthermore, the study showcased the versatility of ACTISM, demonstrating its potential applicability to other automotive systems beyond the IVI, making it a robust tool for safeguarding a wide range of vehicle technologies.
Finally, we outline avenues for future research and development in this area, including automated vulnerability management workflows.
@@ 2 Related Work
Various works have attempted to address security modelling in automotive systems.
@@
However, the framework has several limitations. Firstly, in the threat identification process, the authors did not include physical attack vectors, such as accessing and modifying embedded software in vehicle components, in the STRIDE Data Flow Diagram (DFD). This omission could be because of the assumption that physical attack vectors do not involve flows or exchanges of data. Notwithstanding this assumption, physical attack vectors can be exploited to facilitate data flows, such as "USB Man-in-the-Middle (USB MITM)" attacks, where an attacker intercepts and potentially alters the communication between a USB device and a host system. Consequently, physical attack vectors can be a crucial gap in the overall security design of the vehicle.
Secondly, the CVSS version 2.0 framework adopted in Abuabed et al. [1] is no longer current. The newer versions of CVSS include improvements that deliver more granularity and accuracy in the resulting risk assessment. Lastly, the framework does not adequately address threats in later stages of the life cycle. As automotive system life cycles are typically 10 to 20 years or longer, the vehicle system could become increasingly vulnerable to new types of attacks that were not anticipated during the initial security assessment.
Chlup et al. [2] proposed an automated threat modelling tool for the automotive domain. It combines threat modelling and attack tree analysis using automatically-generated attack trees and attack paths. In addition, it uses anti-patterns to perform in-depth analysis in the threat model. However, the approach does not perform attack feasibility estimation and risk assessment, which are key steps in the TARA workflow.
Other literature relevant to automotive cybersecurity includes a study by Girdhar et al. [5] who used the STRIDE methodology to create a threat model focused on post-accident cyberattack event analysis for CAV.
@@
Additionally, Dang et al. [3] used a STRIDE framework to analyse and classify cybersecurity threats in an autonomous vehicle architecture, including the infotainment system. While the authors ranked the threats using CVSS version 3.0 scores, they did not assess the corresponding impact. The authors focused only on cybersecurity threats originating from outside the vehicle. This means that attacks that involve access to the vehicle’s internal architecture from within the vehicle are excluded in the analysis.
@@
The current landscape of security modelling highlights several limitations. Firstly, there are differing assumptions regarding which vehicle assets fall within the trust boundaries of security models. Secondly, the feasibility and impact of cyber-physical attacks are rarely considered in security modelling. These gaps underscore the need for an iterative and cohesive framework to effectively address the evolving challenges in automotive cybersecurity.
3 Threat-informed Dynamic Security Modelling with ACTISM
We propose ACTISM - a methodology for modelling security threats and attack vectors in a dynamic, iterative manner, specifically designed for automotive systems. We will illustrate the application of ACTISM to an IVI system as a running example and apply ACTISM to a real-world case study, where we demonstrate its ability to dynamically adapt to new threats and improve the security of automotive systems. This case study provides concrete evidence of the framework’s applicability and benefits, distinguishing our research from previous work.
Current security modelling frameworks offer valuable insights but lack adaptability to evolving cyber-physical threats. ACTISM addresses this gap by introducing a framework that continuously updates the cybersecurity posture of automotive systems in response to both cyber and physical threats, attack vectors, and system vulnerabilities.
@@ 3 Threat-informed Dynamic Security Modelling with ACTISM
Unlike static models, ACTISM incorporates threat intelligence through continuous monitoring, which ensures that the security model remains adaptive and effective in mitigating emerging risks. This iterative approach is vital because threat and attack models created in the early stages of system life cycles can become outdated or irrelevant over time as new vulnerabilities emerge. This gap widens, particularly when cyber-physical attacks occur like those targeting keyless entry systems [15], breaching the trust boundary within security models.
ACTISM does not replace current threat and attack modelling frameworks. Instead, it streamlines their adoption and utilisation by cybersecurity practitioners. By integrating various security modelling frameworks into an iterative approach, ACTISM enhances the practicality and efficacy of addressing evolving cybersecurity threats and attacker capabilities. ACTISM comprises eight key steps as shown in Figure 1.
@@
To enhance the practical utility of ACTISM among cybersecurity practitioners, we adopt the STRIDE methodology using a STRIDE-per-element approach, where threats are assessed against each element in the Data Flow Diagram. We choose not to use the STRIDE-per-interaction approach, as research has shown it yields a statistically significant lower level of completeness compared to the STRIDE-per-element approach [13]. Furthermore, we utilise HEAVENS version 2.0 [9] for impact assessment. To our knowledge, HEAVENS is the only framework designed to be fully compliant with the latest automotive cybersecurity standards, such as ISO/SAE 21434 and UNECE Regulation 155.
@@
In addition, we employ CVSS version 3.1 to assess attack feasibility, which allows us to leverage contemporary cybersecurity vulnerability knowledge bases. CVSS version 3.1 provides granular metrics and scoring options, considering various factors that affect the exploitability and impact of a vulnerability. This comprehensive approach ensures that ACTISM remains robust and adaptive to the dynamic nature of cybersecurity threats in automotive systems.
4 Applying ACTISM to an In-Vehicle Infotainment (IVI) System
To illustrate the practical application of ACTISM, it is essential to understand the IVI system and its role in automotive cybersecurity.
@@ 4 Applying ACTISM to an In-Vehicle Infotainment (IVI) System
Legacy head units offered basic features such as radio reception and CD playback, with limited connectivity. In contrast, modern IVI systems are advanced digital platforms that integrate various functionalities beyond audio and navigation. These systems serve as central hubs, providing connectivity to external networks, managing complex data streams, enabling access to vehicle diagnostics, and supporting interactive user interfaces for both drivers and passengers.
While modern IVI systems enhance convenience and operational efficiency, they also expose vehicles to increased cybersecurity risks. These systems are vulnerable to traditional threats, such as remote exploits and data breaches, as well as emerging threats like cyber-physical attacks targeting vehicle functionalities.
The complexity of modern IVI systems significantly expands their attack surface. They integrate multiple wireless communication protocols—Wi-Fi, Bluetooth, and cellular networks—and connect to external services and mobile devices, increasing potential vulnerabilities. Malicious actors can exploit weaknesses in wireless protocols or insecure network connections to gain unauthorised access. Furthermore, advanced functionalities like voice recognition and artificial intelligence add layers of complexity, creating additional attack vectors.
@@
A dynamic approach like ACTISM allows for iterative security modelling throughout the IVI system’s life cycle, ensuring comprehensive protection against evolving threats.

### v3

@@ Abstract
Evolving cybersecurity threats in complex cyber-physical systems pose significant risks to system functionality and safety. This experience report introduces ACTISM (Automotive Consequence-Driven and Threat-Informed Security Modelling), an integrated security modelling framework that enhances the resilience of automotive systems by dynamically updating their cybersecurity posture in response to prevailing and evolving threats, attacker tactics, and their impact on system functionality and safety. ACTISM addresses the existing knowledge gap in static security assessment methodologies by providing a dynamic and iterative framework. We demonstrate the effectiveness of ACTISM by applying it to a real-world example of the Tesla Electric Vehicle’s In-Vehicle Infotainment system, illustrating how the security model can be adapted as new threats emerge. We also report the results of a practitioners’ survey on the usefulness of ACTISM and its future directions. The survey highlights avenues for future research and development in this area, including automated vulnerability management workflows for automotive systems.
@@ 1 Introduction
Automotive systems have evolved into sophisticated cyber-physical systems, packed with cutting-edge technologies. Unlike their predecessors, today’s vehicles are equipped with advanced sensors, including LiDAR (Light Detection and Ranging), radar, and cameras, alongside a variety of connectivity options such as Control Area Networks (CAN), Local Interconnect Networks (LIN), FlexRay, and Ethernet. External connectivity is further enhanced through Bluetooth, Wi-Fi, and GPS, enabling seamless integration with external devices and services. Additionally, the integration of feature-rich In-Vehicle Infotainment (IVI) systems and powerful computing units for autonomous driving continues to push the boundaries of automotive innovation. However, this technological evolution significantly broadens the attack surface, making automotive systems increasingly vulnerable to cybersecurity threats.
Cybersecurity threats and vulnerabilities in the automotive industry are well-documented, with attackers exploiting these weaknesses to compromise vehicle security, access sensitive data, and even take control of critical functions.
Recent literature underscores the growing challenges of defending automotive systems against such threats. For instance, Pham and Xiong [13], in their survey on security attacks and defence techniques for Connected and Autonomous Vehicles (CAVs), highlighted how attackers can sequentially compromise vehicle components. They stressed the importance of securing all system elements to ensure comprehensive protection. While this approach is viable during the early stages of a vehicle’s life cycle, it becomes less effective over time as cybersecurity threats evolve in both complexity and scale. The rapid advancement of automotive technologies further exacerbates this issue, making it impractical to rely solely on initial cybersecurity measures to address emerging risks throughout a vehicle’s operational life.
This is where dynamic threat and attack models prove invaluable. By providing a proactive and adaptive framework, these models enable engineers to continuously analyse and mitigate cybersecurity risks, ensuring systems remain resilient against evolving threats. Additionally, they empower consumers to make informed decisions by evaluating a vehicle’s cybersecurity robustness. However, many existing threat and attack models do not fully address the unique cybersecurity requirements of modern automotive systems. These requirements include:
@@
For instance, threat models developed using Microsoft’s STRIDE (Spoofing, Tampering, Repudiation, Information Disclosure, Denial of Service, and Elevation of Privilege) primarily focus on data flows and information security. As a result, current threat modelling practices for automotive systems often fail to consider the broader consequences of cybersecurity attacks beyond information security [5]. They often overlook critical aspects of how cybersecurity breaches can directly impact automotive functions and safety, raising questions about the reliability of their outcomes when applied to the unique characteristics of automotive environments. Similarly, attack models designed for traditional IT systems do not fully account for the multi-path tactics and techniques attackers use to gain control of a vehicle [2][6]. These techniques often involve a combination of physical and cyber actions, highlighting the need for a more comprehensive and dynamic approach to security modelling.
@@
These findings underscore the importance of tailoring security modelling approaches to automotive systems, ensuring they adequately capture the unique cybersecurity threats these environments face.
@@
Contributions. In this experience report, we introduce ACTISM (Automotive Consequence-Driven and Threat-Informed Security Modelling) — an integrated security modelling framework that enhances the resilience of automotive systems by dynamically updating their cybersecurity posture in response to prevailing and evolving threats, attacker tactics, and their impact on system functionality and safety. This approach addresses the existing knowledge gap in static security assessment methodologies by providing a dynamic and iterative framework. We demonstrate the effectiveness of ACTISM by applying it to an IVI system, showing how the model can be adapted as new threats emerge.
@@
Unlike existing security modelling frameworks, ACTISM integrates multiple methodologies to dynamically update security models with new information gathered through threat intelligence and continuous security monitoring. This ensures that the models remain relevant against evolving attacker tactics, techniques, and procedures (TTPs), even if they were initially developed during the early stages of system design. ACTISM also enables practitioners to apply these security models across both IT and automotive domains by utilising attack paths to accurately represent the multi-path and multi-agent nature of real-world cyber-physical attacks. To enhance the comprehensiveness of its security model, ACTISM applies and adapts the following threat analysis and risk assessment frameworks:
  • STRIDE, a widely-used threat modelling framework developed by Microsoft to characterize security threats in software systems. The acronym STRIDE stands for six types of threats: Spoofing, Tampering, Repudiation, Information Disclosure, Denial of Service, and Elevation of Privilege. In addition to information security, ACTISM reports how physical threats can be modeled using STRIDE in conjunction with other frameworks (Section 4).
@@
  • HEAVENS (i.e., HEAling Vulnerabilities to ENhance Software Security and Safety), a structured and model-based hazard analysis and risk assessment method specifically designed for automotive systems.
@@
  • Attack Tree, a hierarchical diagram representing potential attack paths, used for modeling how an attacker could exploit system vulnerabilities to achieve specific goals.
@@
  • Common Vulnerability Scoring System (CVSS), an open industry standard for assessing the severity of computer system security vulnerabilities. This model is typically used in a static manner, i.e., the severity scores are usually not updated for new or evolving threats. ACTISM shows how CVSS can be applied dynamically (Section 5).
@@
  • Threat Analysis and Risk Assessment (TARA) methodology adopted from the ISO/SAE 21434:2021 ("Road Vehicles - Cybersecurity Engineering") Standard, jointly developed by the International Organisation for Standardisation (ISO) and the Society of Automotive Engineers (SAE), for managing cybersecurity threats and vulnerabilities throughout the vehicle’s life cycle.
@@
While frameworks like STRIDE and CVSS were originally designed for IT security, they can also be applied to physical threats. For instance, tampering with physical components can lead to unauthorised access, data manipulation, or service disruption. ACTISM takes an integrative approach, combining these frameworks to model both cyber and physical attacks. This ensures comprehensive coverage of potential entry points and attack methods, addressing the multi-faceted nature of modern automotive threats.
To demonstrate ACTISM’s practical application, we developed a detailed security model for an In-Vehicle Infotainment (IVI) system. The step-by-step process highlights the integration of threat intelligence and contextual information to accurately characterise and assess cybersecurity risks. Additionally, we illustrate ACTISM’s adaptive capabilities by simulating its response to a real-world vulnerability disclosure affecting the IVI system. This case study underscores ACTISM’s effectiveness in addressing emerging threats and evolving security challenges, while also reporting its versatility for application across other automotive systems beyond IVI.
Finally, we report the results of a practitioners’ survey on the usefulness of ACTISM and its future directions. The survey highlights avenues for future research and development in this area, including automated vulnerability management workflows for automotive systems. These advancements aim to further enhance ACTISM’s ability to safeguard automotive systems against increasingly sophisticated cybersecurity threats.
@@ 2 Related Work
However, the framework has several limitations. Firstly, in the threat identification process, the authors did not include physical attack vectors, such as accessing and modifying embedded software in vehicle components, in the STRIDE Data Flow Diagram (DFD). This omission could be because of the assumption that physical attack vectors do not involve flows or exchanges of data. Notwithstanding this assumption, physical attack vectors can be exploited to facilitate data flows, such as "USB Man-in-the-Middle (USB MITM)" attacks, where an attacker intercepts and potentially alters the communication between a USB device and a host system. Consequently, physical attack vectors can be a crucial gap in the overall security design of the vehicle. Secondly, the CVSS version 2.0 framework adopted in Abuabed et al. [1] lacks the granularity and accuracy that newer versions of CVSS can provide, which impacts the resulting risk assessment. Lastly, the framework does not adequately address threats in later stages of the life cycle. As automotive system life cycles are typically 10 to 20 years or longer, the vehicle system could become increasingly vulnerable to new types of attacks that were not anticipated during the initial security assessment.
@@
Other literature relevant to automotive cybersecurity includes a study by Girdhar et al. [7] who used the STRIDE methodology to create a threat model focused on post-accident cyberattack event analysis for CAV. Chlup et al. [2] proposed an automated threat modelling tool for the automotive domain. It combines threat modelling and attack tree analysis using automatically-generated attack trees and attack paths. In addition, it uses anti-patterns to perform in-depth analysis in the threat model. However, the approach does not perform attack feasibility estimation and risk assessment, which are key steps in the TARA workflow. Dang et al. [3] used a STRIDE framework to analyse and classify cybersecurity threats in an autonomous vehicle architecture, including the infotainment system. While the authors ranked the threats using CVSS version 3.0 scores, they did not assess the corresponding impact. The authors focused only on cybersecurity threats originating from outside the vehicle. This means that attacks that involve access to the vehicle’s internal architecture from within the vehicle are excluded in the analysis.
@@
In summary, the current landscape of security modelling highlights several limitations. Firstly, there are differing assumptions regarding which vehicle assets fall within the trust boundaries of security models. Secondly, the feasibility and impact of cyber-physical attacks are rarely considered in security modelling. These gaps underscore the need for an iterative and cohesive framework to effectively address the evolving challenges in automotive cybersecurity.
3 High-Level Overview of ACTISM
@@
We propose ACTISM, a framework for modelling security threats and attack vectors in a dynamic, iterative manner, specifically designed for automotive systems. Unlike static models, ACTISM incorporates threat intelligence through continuous monitoring, which ensures that the security model remains adaptive and effective in mitigating emerging risks, like those targeting keyless entry systems [16], breaching the trust boundary within security models.
@@
ACTISM is not designed to replace existing threat and attack modelling frameworks. Rather, it simplifies their adoption and application for cybersecurity practitioners. By integrating multiple security modelling frameworks into a cohesive, iterative process, ACTISM significantly improves the practicality and effectiveness of addressing evolving cybersecurity threats and attacker capabilities.
As illustrated in Figure 1, ACTISM consists of eight key steps. These steps integrate multiple other threat and risk assessment frameworks, such as STRIDE, HEAVENS, attack trees, CVSS, and TARA, as outlined in Section 1. The details of each step will be expanded upon in Section 4.
@@
To enhance the practical utility of ACTISM among cybersecurity practitioners, we utilise HEAVENS version 2.0 [11] for impact assessment. To our knowledge, HEAVENS is the only framework designed to be fully compliant with the latest automotive cybersecurity standards, such as ISO/SAE 21434 and UNECE Regulation 155. In addition, we employ CVSS version 3.1 to assess attack feasibility, which allows us to leverage contemporary cybersecurity vulnerability knowledge bases. CVSS version 3.1 provides granular metrics and scoring options, considering various factors that affect the exploitability and impact of a vulnerability. This comprehensive approach ensures that ACTISM remains robust and adaptive to the dynamic nature of cybersecurity threats in automotive systems.
4 Applying ACTISM to an IVI System
To demonstrate ACTISM’s utility, we use an In-Vehicle Infotainment (IVI) system as a running example and apply it to a real-world case study. To appreciate the significance of this application, it is essential to understand the role of IVI systems in modern vehicles and their associated cybersecurity challenges.
@@ 4 Applying ACTISM to an IVI System
Legacy head units offered basic functionalities such as radio reception and CD playback, with minimal connectivity. In contrast, modern IVI systems are sophisticated digital platforms that integrate a wide range of features beyond audio and navigation. They act as central hubs, managing complex data streams, enabling access to vehicle diagnostics, and providing interactive user interfaces for drivers and passengers. Additionally, they connect to external networks, mobile devices, and cloud-based services, significantly enhancing convenience and operational efficiency.
However, these advancements also introduce substantial cybersecurity risks. Modern IVI systems are vulnerable to traditional threats like remote exploits and data breaches, as well as emerging cyber-physical attacks that target vehicle functionalities. The integration of multiple wireless communication protocols such as Wi-Fi, Bluetooth, and cellular networks expands the attack surface, providing malicious actors with opportunities to exploit weaknesses in wireless protocols or insecure network connections. Advanced features like voice recognition and artificial intelligence further increase complexity, creating additional attack vectors.
@@
A dynamic framework like ACTISM is uniquely suited to address these challenges. By enabling iterative security modelling throughout the IVI system’s life cycle, ACTISM ensures comprehensive protection against evolving threats, safeguarding both functionality and safety in modern vehicles.